OpenStudy (lucaz):

how to calculate 1-(1/2)^100

hartnn (hartnn):

why not use a calculator ?

OpenStudy (lucaz):

I need to know how to do it

OpenStudy (lucaz):

my calculator says it's 1

hartnn (hartnn):

ok, whats the original question which brought to this ? anyways, (1/2)^100 = 1/ 2^100 2^100 is quite a large number , so 1/2^100 is very very small number compared to 1 approximately it will = 1

OpenStudy (lucaz):

hmm, I thought so, it's from a geometric series

hartnn (hartnn):

like sum of 1st hundred numbers with common ratio = 1/2 ? and 1st term = 1/2 right ?

OpenStudy (lucaz):

the rest is ok, I was just wanted to know if there was some trick to calculate the 1/2^100.. thanks!

hartnn (hartnn):

ok, welcome ^_^